Can you explain the paper Good and Bad Uncertainty: Macroeconomic and Financial Market Implications (Gill Segal et. al) with math examples?

1
Expert's answer
2021-09-21T10:59:23-0400

Consider an economy with a continuum of firms with productivity "\theta^i \backsim N(\bar{\theta},\sigma^2_\theta)."

Assume linear technologies: "\frac{dB^i_t}{B^i_t}=\theta^idt+\sigma dW_t+\sigma_1dW_{i,t}"

Aggregate capital "B_t=\int B^i_tdt"

The law of large numbers "\implies\frac{dB_t}{B_t}=(\bar{\theta}+\frac{1}{2}\sigma^2_\theta t)dt+\sigma dW_{t}"

 higher dispersion of productivity increase the drift rate of aggregate capital (and hence, investments, consumption etc).

Moreover, aggregate price is increasing in "\u03c3_\u03b8" . Given a log-normal SDF and assuming only one dividend

"D_T = B_T" at T we have

"\frac{M_t}{B_t}=e^{(\bar{\theta}-r-\sigma \sigma_M)(T-t)+\frac{1}{2}\sigma^2_\theta(T-t)^2}"

In this model:

"\u03c3_\u03b8 =" Good uncertainty

"\u03c3, \u03c3_M =" Bad uncertainty